# A Malliavin-Gamma calculus approach to Score Based Diffusion Generative models for random fields

### v1

@@ Abstract.
We warmly thank Giovanni Conforti, Lorenzo Dello Schiavo and Domenico Marinucci for many useful discussions. While this work was written GG was associated to INdAM (Istituto Nazionale di Alta Matematica “Francesco Severi”) and the group GNAMPA. GG is supported by the PRIN project GRAFIA (CUP: E53D23005530006) and MUR Departement of Excellence Programme 2023-2027 MatMod@Tov (CUP: E83C23000330006).
@@ 1. Introduction
The aim of generative modelling can be summarized as creating data from noise [42]. In practice this means aiming at generating new data from an unknown distribution $\mu_{\mathrm{data}}$, having access only to a bunch of samples. Among all existing methods, the most promising ones (and in practice also the most employed) are Score-based diffusion Generative Models (SGMs) [42, 17, 38, 11]. The core mechanism of SGMs involves two primary processes: a forward diffusion process and a reverse generation process. In the forward process, data is progressively corrupted by adding noise, effectively transforming it into a simple, known distribution (e.g., Gaussian noise). This process can be mathematically described by a stochastic differential equation (SDE) that models the gradual addition of noise over time. The reverse process aims to recover the original data by learning to denoise the noisy data step by step. This is achieved by estimating the score function, which in the Euclidean setting corresponds to the gradient of the log-density of the forward noising process. By leveraging this score function, the model can iteratively refine noisy samples back into data samples that resemble the original data distribution. This reverse process is also governed by an SDE, which, when solved numerically, facilitates the generation of new data samples from pure noise. To be more precise, in the classical Euclidean setting $\mathbb{R}^{d}$ the idea is to consider a noising forward diffusion process started in our data samples (distributed according to the unknwon $\mu_{\mathrm{data}}$), that converges to a known (and easy-to-sample) noising target distribution. A standard choice would be choosing a target Gaussian distribution $\mathcal{N}(0,\lambda)$, whose density equals $\mathfrak{m}\propto\exp(-\nicefrac{{|x|^{2}}}{{2\lambda}})\mathrm{d}x$, which means considering as forward noising process the Ornstein–Uhlenbeck process
@@
(1.2) $\begin{cases}\mathrm{d}\overleftarrow{X}_{t}=-\frac{1}{2}\overleftarrow{X}_{t}\mathrm{d}t+\lambda\nabla\log\frac{d\overrightarrow{\mathrm{P}}_{T-t}}{d\mathfrak{m}}(\overleftarrow{X}_{t})\mathrm{d}t+\sqrt{\lambda}\mathrm{d}B_{t}\\ \overleftarrow{X}_{T}\sim\mu_{\mathrm{data}}\,,\end{cases}$
@@
The core idea behind this paper is using a more abstract point of view, based on $\Gamma$-calculus and Dirichlet forms, in order to define an appropriate forward noising dynamics. In doing so, we unveil the role played by Malliavin calculus, Malliavin derivatives, and Wiener chaoses, and we combine it with abstract time-reversal formulas for identifying the backward denoising process. Despite we restrict ourselves here to a Gaussian setting (i.e., the noising dynamics converges to a Gaussian measure), we believe that what we present below might be used as a guideline for engineering infinite-dimensional SGMs with different target noises, which one would like to consider taking into account the properties of $\mu_{\mathrm{data}}$ (e.g., sparsity). Lastly, we employ this general approach in order to get entropic convergence bounds in $\mathrm{KL}$ divergence.
@@ 1.2. Main contributions
Lastly, we conclude the paper with Section 5 where we consider as state space the space of spherical random fields and where we consider as noisy Gaussian measure the law of a Whittle-Matérn random field.
@@ 2. Noising dynamics in Hilbertian setting
The state space of our Markov forward noising diffusion is the Hilbert space $(\mathrm{H},\langle\cdot,\cdot\rangle_{\mathrm{H}})$ (as running example we will think of $\mathrm{H}={\mathrm{L}^{2}({\mathbb{S}^{2}})}$ equipped with its scalar product, induced by the volume measure $\mathrm{vol}$). We identify the dual of $\mathrm{H}$ (namely, $\mathrm{H}^{\star}$) with the Hilbert space itself via Riesz Theorem and denote with $\varphi^{\star}\in\mathrm{H}^{\star}$ the bounded linear functional $u\mapsto\langle\varphi,u\rangle_{H}$ on $\mathrm{H}$, associated to the element $\varphi\in\mathrm{H}$. Next, we should introduce our noising Markovian evolution. This can be done by firstly considering a Markov triple structure $(E,\mathfrak{m},\Gamma)$, that is a measure space $(E,\mathfrak{m})$ and a carré du champ operatpr $\Gamma$, defined on some appropriate set of functions. The measure space considered here is the Hilbert space $\mathrm{H}$ equipped with the Gaussian measure $\mathfrak{m}$ with symmetric positive definite covariance operator $C$. We assume $C$ to be of trace class, that is we assume $C$ to have the spectral decomposition $CY_{\ell}=C_{\ell}Y_{\ell}$, with $(Y_{\ell})_{\ell\geq 0}$ being an orthonormal base of $\mathrm{H}$, with $C_{\ell}>0$ and $\sum_{\ell}C_{\ell}<\infty$. As running example in $\mathrm{H}={\mathrm{L}^{2}({\mathbb{S}^{2}})}$, the reader could consider $\mathfrak{m}$ being the law of a Whittle-Matérn spherical random field, $(Y_{\ell})_{\ell\geq 0}$ being the spherical harmonics and $(C_{\ell})_{\ell\geq 0}$ being the angular power spectrum of the Whittle-Matérn field, cf. Section 5 below.
@@
$\mathcal{PC}\coloneqq\biggl{\{}p=F(I(f_{1}),\,\dots,\,I(f_{k}))\quad\text{ forsome }k\geq 1,\,F\in\mathcal{C}^{\infty}_{\mathrm{Pol}}(\mathbb{R}^{k}),\,\text{ and }f_{i}\in\mathcal{H}_{\mathfrak{m}}\quad\forall\,i=1,\dots,k\biggr{\}}\,,$
@@
$\langle h,k\rangle_{\mathcal{H}_{\mathfrak{m}}}\coloneqq\langle C^{-\nicefrac{{1}}{{2}}}h,C^{-\nicefrac{{1}}{{2}}}k\rangle_{\mathrm{H}}=\sum_{k,l}\frac{\langle h,Y_{\ell}\rangle_{H}\langle k,Y_{\ell}\rangle_{H}}{C_{\ell}}\,.$
@@
and for notations’ sake we will also summarize the above display by simply writing $p=F(I(\underline{f}))=F\circ I(\underline{f})$. Clearly we have $\mathcal{R}_{\mathfrak{m}}\subset\mathcal{PC}\subset\mathrm{L}^{2}(\mathrm{H},\mathfrak{m})$. Moreover, $\mathcal{PC}$ is an algebra of nowhere-vanishing functions that separates points which is dense in $\mathrm{L}^{2}(\mathrm{H},\mathfrak{m})$, the latter property being a straightforward consequence of the Wiener chaos decomposition [29, Theorem 1.1.1].
@@
(2.1) ¯$\bar{\nabla}p=\sum_{i=1}^{k}(\partial_{i}F)(I(f_{1}),\dots,I(f_{k}))\,f_{i}\,\in\mathcal{H}_{\mathfrak{m}}\,.$
Notice that for any $f\in\mathcal{H}_{\mathfrak{m}}$ we immediately have ¯$\bar{\nabla}I(f)=f\in\mathcal{H}_{\mathfrak{m}}$ and recall that the gradient operator $\bar{\nabla}$ here defined is closable from $\mathrm{L}^{p}(\mathrm{H},\mathfrak{m})$ to $\mathrm{L}^{p}(\mathrm{H};\mathcal{H}_{\mathfrak{m}})$ for any $p\geq 1$ [28, Proposition 2.3.4]. Then we may finally define the carré du champ operator $\Gamma$ on $\mathcal{PC}$ as
@@
$\mathcal{E}(F(I(\underline{f})),G(I(\underline{g})))=\frac{1}{2}\int_{\mathrm{H}}\Gamma(F(I(\underline{f})),G(I(\underline{g})))\mathrm{d}\mathfrak{m}=\frac{1}{2}\mathbb{E}\biggl{[}\left\langle\bar{\nabla}(F\circ I(\underline{f}))(\Phi),\bar{\nabla}(G\circ I(\underline{g}))(\Phi)\,\right\rangle_{\mathcal{H}_{\mathfrak{m}}}\biggr{]}$
@@
where in the last step we have relied on the integration by parts formula for Malliavin calculus [29, Lemma 1.2.1].
Therefore, we may define the operator $L$, at least on the reproducing kernel Hilbert space $\mathcal{R}_{\mathfrak{m}}$, as $LI(w)=-I(w)/2$. In order to extend it to the whole algebra $\mathcal{PC}$ it is then enough imposing the chain rule/diffusion property (inherited from $\Gamma$)
@@
Let us notice that the integration by parts step applied in (2.3) can be equivalently deduced from the Cameron-Martin Theorem. We sketch here the formal proof since it highlights the importance of the Cameron-Martin space $\mathcal{H}_{\mathfrak{m}}$. The main idea is the simple observation (see [29, Definition 1.2.1] and following discussion) that
¯$\left\langle\bar{\nabla}(G\circ I(\underline{g}))(\Phi),\,w\,\right\rangle_{\mathcal{H}_{\mathfrak{m}}}=\lim_{t\downarrow 0}\frac{1}{t}[G\circ I(\underline{g})(\Phi+tw)-G\circ I(\underline{g})(\Phi)]\,.$
@@
$=\frac{1}{2}\int(G\circ$ $I(\underline{g}))\,\biggl{(}\mathrm{d}_{t}|_{t=0}\frac{\mathrm{d}(e^{tw}.)_{\#}\mathfrak{m}}{\mathrm{d}\mathfrak{m}}\biggr{)}\mathrm{d}\mathfrak{m}=\frac{1}{2}\int(G\circ I(\underline{g}))\,I(w)\,\mathrm{d}\mathfrak{m}=\frac{1}{2}\mathbb{E}[(G\circ I(\underline{g}))\,I(w)]\,.$
@@
$\mathcal{D}(L)\coloneqq\{p\in\mathcal{D}(\mathcal{E})\quad\colon\quad\exists\,C(p)>0\text{ s.t. }\mathcal{E}(p,q)\leq C(p)\|q\|_{\mathrm{L}^{2}(\mathrm{H},\mathfrak{m})}\quad\forall\,q\in\mathcal{D}(\mathcal{E})\}\,.$
@@
Moreover, it is positive preserving and sub-Markov (cf [6, Section 1.3.5]). Actually, this semigroup is well-known and it corresponds to OU process on $\mathrm{H}$ (see for instance [29, Section 1.4]). To see this, is enough noticing that our generator and the OU generator coincide on our cylindrical algebra $\mathcal{PC}$ [29, Proposition 1.4.4], that $\mathcal{D}(L)=\mathbb{D}^{2,2}$ (the completion of $\mathcal{PC}$ under the norm $\|\phi\|_{2,2}=(\|\phi\|^{2}_{\mathrm{L}^{2}(\mathrm{H},\mathfrak{m})}+\|\bar{\nabla}\phi\|^{2}_{\mathcal{H}_{\mathfrak{m}}}+\|\bar{\nabla}^{2}\phi\|^{2}_{\mathcal{H}_{\mathfrak{m}}^{\otimes 2}})^{\nicefrac{{1}}{{2}}}$) and that $L$ is essentially self-adjoint, i.e., that $\mathcal{PC}$ is a core for $L$ (meaning that $\mathcal{PC}$ is dense in $\mathcal{D}(L)$ in the graph norm $\|\cdot\|_{2,2}$) [29, Corollary 1.5.1]. For the construction of this Dirichlet structure we refer the reader also to [27, Chapter II Section 3]. Lastly, owing to [29, Section 1.4] we may express $L$ and the corresponding semigroup $(P_{t})_{t\geq 0}$ in terms of Wiener chaoses. In view of that, let us consider the Wiener chaos decomposition
@@
and $\mathcal{D}(L)=\{u\in\mathrm{L}^{2}(\mathrm{H},\mathfrak{m})\,\colon\,\sum_{n\geq 0}n^{2}\,\left\|\Pi_{n}u\right\|_{\mathrm{L}^{2}(\mathrm{H},\mathfrak{m})}^{2}<+\infty\}$. Finally, owing to Mehler’s formula [29, Equation 1.67] the semigroup $P_{t}$ may be extended to $\mathrm{L}^{1}(\mathrm{H},\mathfrak{m})$.
@@
where $(W^{\mathfrak{m}}_{t})_{t\geq 0}$ is an $(\mathcal{F}_{t})_{t}$-Brownian motion on $\mathrm{H}$ starting at zero with covariance $\langle\cdot,\cdot\rangle_{\mathcal{H}_{\mathfrak{m}}}$ (cf. [5, Theorem 6.10 and Remark 6.8-(ii)] or [27, Theorem II.3.11]) and for any $t\geq 0$, for any $u\in\mathrm{L}^{2}(\mathrm{H},\mathfrak{m})$ it holds $P_{t}u(x)=\mathbb{E}[u(X_{t})|X_{0}=x]$ for $\mathfrak{m}$-a.e. $x\in\mathrm{H}$. Moreover, from Fukushima decomposition [5, Theorem 4.3 and Remark 4.4-(ii)] it follows that $\mathrm{P}$ solves the martingale problem $MP(L,\mathcal{D}(L))$, i.e., that for any $u\in\mathcal{D}(L)$ the process
@@
In the forthcoming discussion we are going to assume that $\mathrm{KL}(\mu_{\mathrm{data}}|\mathfrak{m})$ is finite. This guarantees smoothing effects of the semigroup $(P_{t})_{t\geq 0}$. Indeed, if we set $\mu_{t}=(X_{t})_{\#}\overrightarrow{\mathrm{P}}=\mu_{\mathrm{data}}P_{t}$ and $\rho_{t}\coloneqq\nicefrac{{\mathrm{d}\mu_{t}}}{{\mathrm{d}\mathfrak{m}}}=P_{t}\rho_{0}$, then the former can be seen as the gradient flow on $\mathcal{P}_{2}(\mathrm{H})$ of the relative entropy functional [2, Section 10.4.8] which implies Sobolev regularity for $\log\rho_{t}$ [2, Theorem 11.2.12]. More precisely, we may conclude that for a.e. $t>0$ it holds ¯$\int\nicefrac{{\|\bar{\nabla}\rho_{t}\|^{2}_{\mathcal{H}_{\mathfrak{m}}}}}{{\rho_{t}}}\mathrm{d}\mathfrak{m}<+\infty$, i.e., that $\rho_{t}$ has finite Fisher information (w.r.t. $\mathfrak{m}$). Notice that this further implies that $\sqrt{\rho_{t}}\,\in\mathcal{D}(\mathcal{E})$ and $\rho_{t}\,\in\mathcal{D}(\mathcal{E})$ (via Poincaré inequality [7, Theorem 5.5.1]). Finally, our convergence bounds will further require ¯$\left\|\bar{\nabla}\sqrt{\rho_{0}}\right\|_{\mathcal{H}_{\mathfrak{m}}}$ to be finite, i.e., that $\sqrt{\rho_{0}}\in\mathcal{D}(\mathcal{E})$.

### v2

@@ Abstract.
We warmly thank Giovanni Conforti, Lorenzo Dello Schiavo and Domenico Marinucci for many useful discussions. While this work was written $G$ was associated to INdAM (Istituto Nazionale di Alta Matematica “Francesco Severi”) and the group GNAMPA. $G$ is supported by the PRIN project GRAFIA (CUP: E53D23005530006) and MUR Departement of Excellence Programme 2023-2027 MatMod@Tov (CUP: E83C23000330006).
@@ 1. Introduction
The aim of generative modelling can be summarized as creating data from noise [45]. In practice this means aiming at generating new data from an unknown distribution $\mu_{\mathrm{data}}$, having access only to a bunch of samples. Among all existing methods, the most promising ones (and in practice also the most employed) are Score-based diffusion Generative Models (SGMs) [45, 20, 42, 12]. The core mechanism of SGMs involves two primary processes: a forward diffusion process and a reverse generation process. In the forward process, data is progressively corrupted by adding noise, effectively transforming it into a simple, known distribution (e.g., Gaussian noise). This process can be mathematically described by a stochastic differential equation (SDE) that models the gradual addition of noise over time. The reverse process aims to recover the original data by learning to denoise the noisy data step by step. This is achieved by estimating the score function, which in the Euclidean setting corresponds to the gradient of the log-density of the forward noising process. By leveraging this score function, the model can iteratively refine noisy samples back into data samples that resemble the original data distribution. This reverse process is also governed by an SDE, which, when solved numerically, facilitates the generation of new data samples from pure noise. To be more precise, in the classical Euclidean setting $\mathbb{R}^{d}$ the idea is to consider a noising forward diffusion process started in our data samples (distributed according to the unknown $\mu_{\mathrm{data}}$), that converges to a known (and easy-to-sample) noising target distribution. A standard choice would be choosing a target Gaussian distribution $\mathcal{N}(0,\lambda)$, whose density equals $\mathfrak{m}\propto\exp(-\nicefrac{{|x|^{2}}}{{2\lambda}})\mathrm{d}x$, which means considering as forward noising process the Ornstein–Uhlenbeck process
@@
[…]
@@
The core idea behind this paper is using a more abstract point of view, based on $\Gamma$-calculus and Dirichlet forms, in order to define an appropriate forward noising dynamics. In doing so, we unveil the role played by Malliavin calculus, Malliavin derivatives, and Wiener chaoses, and we combine it with abstract time-reversal formulas for identifying the backward denoising process. Despite we restrict ourselves here to a Gaussian setting (i.e., the noising dynamics converges to a Gaussian measure), we believe that what we present below might be used as a guideline for engineering infinite-dimensional SGMs with different target noises, which one would like to consider taking into account the properties of $\mu_{\mathrm{data}}$ (e.g., sparsity). Then, we employ this general approach in order to get entropic convergence bounds in $\mathrm{KL}$ divergence.
Lastly, let us simply mention here that the main difficulty for going beyond the pure Gaussian setting in the infinite-dimensional setting might be the lack of explicit known identities for the corresponding noising stochastic process associated to the abstract Dirichlet form. This would prevent any practical application. Moreover, in order to compute the score functional we rely on some explicit properties of the noising process. Therefore, any non-Gaussian generalization that builds upon the present work would still require the explicit knowledge of the noising process. We will address more in depth this important generalization in a follow-up work.
@@ 1.2. Main contributions
Lastly, we conclude the paper with Section 4.1 where we consider as state space the space of spherical random fields and where we consider as noisy Gaussian measure the law of a Whittle-Matérn random field.
@@
While this paper was under review, two recent contributions [30, 31] have appeared that are closely related to ours. In the first, the authors combine Malliavin derivatives and divergences with a new Bismut-type formula in order to deduce exact, closed-form, expression for the score function for a broad class of non-linear diffusion generative models in finite-dimensional settings. In the latter, the authors manage to extend their work from finite- to infinite-dimensional diffusion models, using techniques for differentiating Hilbert-valued processes together with infinite-dimensional extensions of the Bismut–Elworthy–Li formula. In particular, they obtain closed-form expressions for the score function as well in broadly general settings.
@@ 2. Noising dynamics in Hilbertian setting
The state space of our Markov forward noising diffusion is the Hilbert space $(\mathrm{H},\langle\cdot,\cdot\rangle_{\mathrm{H}})$ (as running example we will think of $\mathrm{H}={\mathrm{L}^{2}({\mathbb{S}^{2}})}$ equipped with its scalar product, induced by the volume measure $\mathrm{vol}$). We identify the dual of $\mathrm{H}$ (namely, $\mathrm{H}^{\star}$) with the Hilbert space itself via Riesz Theorem and denote with $\varphi^{\star}\in\mathrm{H}^{\star}$ the bounded linear functional $u\mapsto\langle\varphi,u\rangle_{H}$ on $\mathrm{H}$, associated to the element $\varphi\in\mathrm{H}$. Next, we should introduce our noising Markovian evolution. This can be done by firstly considering a Markov triple structure $(E,\mathfrak{m},\Gamma)$, that is a measure space $(E,\mathfrak{m})$ and a carré du champ operatpr $\Gamma$, defined on some appropriate set of functions. The measure space considered here is the Hilbert space $\mathrm{H}$ equipped with the Gaussian measure $\mathfrak{m}$ with symmetric positive definite covariance operator $C$. We assume $C$ to be of trace class, that is we assume $C$ to have the spectral decomposition $CY_{\ell}=C_{\ell}Y_{\ell}$, with $(Y_{\ell})_{\ell\geq 0}$ being an orthonormal base of $\mathrm{H}$, with $C_{\ell}>0$ and $\sum_{\ell}C_{\ell}<\infty$. As running example in $\mathrm{H}={\mathrm{L}^{2}({\mathbb{S}^{2}})}$, the reader could consider $\mathfrak{m}$ being the law of a Whittle-Matérn spherical random field, $(Y_{\ell})_{\ell\geq 0}$ being the spherical harmonics and $(C_{\ell})_{\ell\geq 0}$ being the angular power spectrum of the Whittle-Matérn field, cf. Section 4.1 below.
@@
$\mathcal{PC}\coloneqq\biggl\{p=F(I(f_{1}),\,\dots,\,I(f_{k}))\quad\colon\quad\begin{aligned} \,&\text{for some }k\geq 1,\,F\in\mathcal{C}^{\infty}_{\mathrm{Pol}}(\mathbb{R}^{k})\\ \,&\text{and }f_{i}\in\mathcal{H}_{\mathfrak{m}}\quad\forall\,i=1,\dots,k\end{aligned}\biggr\}\,$,
@@
$\langle h,k\rangle_{\mathcal{H}_{\mathfrak{m}}}\coloneqq\langle C^{-\nicefrac{{1}}{{2}}}h,C^{-\nicefrac{{1}}{{2}}}k\rangle_{\mathrm{H}}=\sum_{\ell\geq 0}\frac{\langle h,Y_{\ell}\rangle_{H}\langle k,Y_{\ell}\rangle_{H}}{C_{\ell}}\,.$
@@
and for notations’ sake we will also summarize the above display by simply writing $p=F(I(\underline{f}))=F\circ I(\underline{f})$. Clearly we have $\mathcal{R}_{\mathfrak{m}}\subset\mathcal{PC}\subset\mathrm{L}^{2}(\mathrm{H},\mathfrak{m})$. Moreover, $\mathcal{PC}$ is an algebra of nowhere-vanishing functions that separates points which is dense in $\mathrm{L}^{2}(\mathrm{H},\mathfrak{m})$, the latter property being a straightforward consequence of the Wiener chaos decomposition [34, Theorem 1.1.1].
@@
(2.1) $\bar{\nabla}p=\sum_{i=1}^{k}(\partial_{i}F)(I(f_{1}),\dots,I(f_{k}))\,f_{i}\,\in\mathcal{H}_{\mathfrak{m}}\,.$
Notice that for any $f\in\mathcal{H}_{\mathfrak{m}}$ we immediately have $\bar{\nabla}I(f)=f\in\mathcal{H}_{\mathfrak{m}}$ and recall that the gradient operator $\bar{\nabla}$ here defined is closable from $\mathrm{L}^{p}(\mathrm{H},\mathfrak{m})$ to $\mathrm{L}^{p}(\mathrm{H};\mathcal{H}_{\mathfrak{m}})$ for any $p\geq 1$ [33, Proposition 2.3.4]. Then we may finally define the carré du champ operator $\Gamma$ on $\mathcal{PC}$ as
@@
$\mathcal{E}(F(I(\underline{f})),G(I(\underline{g})))=$ $\,\frac{1}{2}\int_{\mathrm{H}}\Gamma(F(I(\underline{f})),G(I(\underline{g})))\mathrm{d}\mathfrak{m}$
$=$ $\,\frac{1}{2}\mathbb{E}\biggl[\left\langle\bar{\nabla}(F\circ I(\underline{f}))(\Phi),\bar{\nabla}(G\circ I(\underline{g}))(\Phi)\,\right\rangle_{\mathcal{H}_{\mathfrak{m}}}\biggr]$
@@
where in the last step we have relied on the integration by parts formula for Malliavin calculus [34, Lemma 1.2.1].
Therefore, we may define the operator $L$, at least on the reproducing kernel Hilbert space $\mathcal{R}_{\mathfrak{m}}$, as LI(w)=I(w)/2LI(w)=-I(w)/2. In order to extend it to the whole algebra $\mathcal{PC}$ it is then enough imposing the chain rule/diffusion property (inherited from $\Gamma$)
@@
Let us notice that the integration by parts step applied in (2.3) can be equivalently deduced from the Cameron-Martin Theorem. We sketch here the formal proof since it highlights the importance of the Cameron-Martin space $\mathcal{H}_{\mathfrak{m}}$. The main idea is the simple observation (see [34, Definition 1.2.1] and following discussion) that
$\left\langle\bar{\nabla}(G\circ I(\underline{g}))(\Phi),\,w\,\right\rangle_{\mathcal{H}_{\mathfrak{m}}}=\lim_{t\downarrow 0}\frac{1}{t}[G\circ I(\underline{g})(\Phi+tw)-G\circ I(\underline{g})(\Phi)]\,.$
@@
$=$ $\,\frac{1}{2}\int(G\circ I(\underline{g}))\,\biggl(\mathrm{d}_{t}|_{t=0}\frac{\mathrm{d}(e^{tw}.)_{\#}\mathfrak{m}}{\mathrm{d}\mathfrak{m}}\biggr)\mathrm{d}\mathfrak{m}$
$=$ $\,\frac{1}{2}\int(G\circ I(\underline{g}))\,I(w)\,\mathrm{d}\mathfrak{m}=\frac{1}{2}\mathbb{E}[(G\circ I(\underline{g}))\,I(w)]\,.$
@@
$\mathcal{D}(L)\coloneqq\{p\in\mathcal{D}(\mathcal{E})\;\colon\;\exists\,C(p)>0\text{ s.t. }\mathcal{E}(p,q)\leq C(p)\|q\|_{\mathrm{L}^{2}(\mathrm{H},\mathfrak{m})}\;\forall\,q\in\mathcal{D}(\mathcal{E})\}\,.$
@@
Moreover, it is positive preserving and sub-Markov (cf [6, Section 1.3.5]). Actually, this semigroup is well-known and it corresponds to OU process on $\mathrm{H}$ (see for instance [34, Section 1.4]). To see this, is enough noticing that our generator and the OU generator coincide on our cylindrical algebra $\mathcal{PC}$ [34, Proposition 1.4.4], that $\mathcal{D}(L)=\mathbb{D}^{2,2}$ (the completion of $\mathcal{PC}$ under the norm $\|\phi\|_{2,2}=(\|\phi\|^{2}_{\mathrm{L}^{2}(\mathrm{H},\mathfrak{m})}+\|\bar{\nabla}\phi\|^{2}_{\mathcal{H}_{\mathfrak{m}}}+\|\bar{\nabla}^{2}\phi\|^{2}_{\mathcal{H}_{\mathfrak{m}}^{\otimes 2}})^{\nicefrac{{1}}{{2}}}$) and that $L$ is essentially self-adjoint, i.e., that $\mathcal{PC}$ is a core for $L$ (meaning that $\mathcal{PC}$ is dense in $\mathcal{D}(L)$ in the graph norm $\|\cdot\|_{2,2}$) [34, Corollary 1.5.1]. For the construction of this Dirichlet structure we refer the reader also to [32, Chapter II Section 3]. Lastly, owing to [34, Section 1.4] we may express $L$ and the corresponding semigroup $(P_{t})_{t\geq 0}$ in terms of Wiener chaoses. In view of that, let us consider the Wiener chaos decomposition
@@
and $\mathcal{D}(L)=\{u\in\mathrm{L}^{2}(\mathrm{H},\mathfrak{m})\,\colon\,\sum_{n\geq 0}n^{2}\,\left\|\Pi_{n}u\right\|_{\mathrm{L}^{2}(\mathrm{H},\mathfrak{m})}^{2}<+\infty\}$. Finally, owing to Mehler’s formula [34, Equation 1.67] the semigroup $P_{t}$ may be extended to $\mathrm{L}^{1}(\mathrm{H},\mathfrak{m})$.
@@
where $(W^{\mathfrak{m}}_{t})_{t\geq 0}$ is an $(\mathcal{F}_{t})_{t}$-Brownian motion on $\mathrm{H}$ starting at zero with covariance $\langle\cdot,\cdot\rangle_{\mathcal{H}_{\mathfrak{m}}}$ (cf. [5, Theorem 6.10 and Remark 6.8-(ii)] or [32, Theorem II.3.11]) and for any $t\geq 0$, for any $u\in\mathrm{L}^{2}(\mathrm{H},\mathfrak{m})$ it holds $P_{t}u(x)=\mathbb{E}[u(X_{t})|X_{0}=x]$ for $\mathfrak{m}$-a.e. $x\in\mathrm{H}$. Moreover, from Fukushima decomposition [5, Theorem 4.3 and Remark 4.4-(ii)] it follows that $\mathrm{P}$ solves the martingale problem $MP(L,\mathcal{D}(L))$, i.e., that for any $u\in\mathcal{D}(L)$ the process
@@
In the forthcoming discussion we are going to assume that $\mathrm{KL}(\mu_{\mathrm{data}}|\mathfrak{m})$ is finite. This guarantees smoothing effects of the semigroup $(P_{t})_{t\geq 0}$. Indeed, if we set $\mu_{t}=(X_{t})_{\#}\overrightarrow{\mathrm{P}}=\mu_{\mathrm{data}}P_{t}$ and $\rho_{t}\coloneqq\nicefrac{{\mathrm{d}\mu_{t}}}{{\mathrm{d}\mathfrak{m}}}=P_{t}\rho_{0}$, then the former can be seen as the gradient flow on $\mathcal{P}_{2}(\mathrm{H})$ of the relative entropy functional [2, Section 10.4.8] which implies Sobolev regularity for $\log\rho_{t}$ [2, Theorem 11.2.12]. More precisely, we may conclude that for a.e. $t>0$ it holds $\int\nicefrac{{\|\bar{\nabla}\rho_{t}\|^{2}_{\mathcal{H}_{\mathfrak{m}}}}}{{\rho_{t}}}\mathrm{d}\mathfrak{m}<+\infty$, i.e., that $\rho_{t}$ has finite Fisher information (w.r.t. $\mathfrak{m}$). Notice that this further implies that $\sqrt{\rho_{t}}\,\in\mathcal{D}(\mathcal{E})$ and $\rho_{t}\,\in\mathcal{D}(\mathcal{E})$ (via Poincaré inequality [7, Theorem 5.5.1]). Finally, our convergence bounds will further require $\left\|\bar{\nabla}\sqrt{\rho_{0}}\right\|_{\mathcal{H}_{\mathfrak{m}}}$ to be finite, i.e., that $\sqrt{\rho_{0}}\in\mathcal{D}(\mathcal{E})$.